Circular No. 5863

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1993 M. Della Valle, European Southern Observatory; and S. Cristiani and

C. Lissandrini, Dipartimento di Astronomia, Universitá di Padova, report: “Analysis of spectrograms (range 370-900 nm, resolution ∼ 0.8-3 nm), obtained on Sept. 16.97 UT with the ESO/MPI 2.2-m telescope (+ EFOSC2) at La Silla, confirms this object (cf. IAUC 5862) as being a nova, caught a few days past maximum. The spectrum shows emission lines of H, Fe II (mult. 42, 49, 55, 73, 74), Na I (589 nm), and O I (mult. 1, 4, 34). The emission line of Na I is flanked by a clear, blue-shifted P-Cyg profile. The expansion velocity derived from the minimum of the P-Cyg absorption and the FWHM of Hα is ∼ 2400 km/s.”

P. M. Kilmartin reports the following precise position from an exposure obtained by A. C. Gilmore at Mt. John University Observatory on Sept. 17.338 UT: α = 18^{h}09^{m}38^{s}.40, δ = −29^{o}29′54′′.0 (equinox 1950.0). Gilmore also provides the following photometry, obtained in marginal conditions with a 0.6-m f/16 Cassegrain telescope (+ Johnson-Cousins UBVR_{c}I_{c} filters): Sept. 17.414, V = 8.59 ± 0.00, U–B = –0.26 ± 0.05, B–V = +0.33 ± 0.00, V–R = +0.42 ± 0.01, V–I = +1.00 ± 0.12; 17.434, 8.57 ± 0.02, –0.30 ± 0.01, +0.35 ± 0.01, +0.40 ± 0.01, +0.96 ± 0.04.

GRO J0422+32 G. Beskin, S. Kajsin, S. Neizvestny, I. Panferova, V. Plokhotnichenko,

A. Tikhonov, A. Ugrumov, and A. Zhuravkov, Special Astrophysical Observatory (SAO); and C. Bartolini, G. Cosentino, A. Guarnieri, and A. Piccioni, University of Bologna, communicate: “Photoelectric and spectroscopic observations from Sept. 9 to 12, obtained at the 6-m SAO telescope (+ Mania complex), confirm that GRO J0422+32 is in a very low state, as indicated by the following Johnson photometry: Sept. 10.02 UT, V = 18.17 ± 0.07, B = 19.32 ± 0.30, R = 17.50 ± 0.25; 10.94, 18.19 ± 0.08, 19.67 ± 0.24, –; 12.03, 18.99 ± 0.10, 19.99 ± 0.20, 18.25 ± 0.20 (1-m SAO telescope). Five spectrograms show only a continuum in the range 400-500 nm, with weak Hα in the range 500-600 nm. Perhaps the rapid fading reported by Filippenko and Matheson (IAUC 5860) is an indication that the drop observed by Bartolini et al. (IAUC 5690) is recurrent.”

1993 September 17 Daniel W. E. Green